OpenStudy (anonymous):

Hello, I have a math question that has me stumped, could any give me the way to figure it out? Tina is twice as old as her cousin Sally. Their ages when added are 21years. How many years old is each girl? Thanks

hero (hero):

t + s = 21 t = 2s Solve the system

OpenStudy (anonymous):

Frame a equation as follows : Let the age of Sally be x ,then the age of Tina is 2x.Thus according to the question : 2x + x = 21 .

hero (hero):

2s + s = 21 3s = 21 s = 7 t = 14

OpenStudy (anonymous):

Thank you so much, I got it now how do I explain this to an 8yr old so she can get it too!

OpenStudy (anonymous):

If one is twice as old as the other, one-third of their total age is half the age of one and the age of the other.